OpenStudy (anonymous):

Hello there I need help with this problem OK so if 6/y + 9/2 y =40 then y =?

OpenStudy (anonymous):

[Y(6/y+9/2y)=40y] 6+9 =40y 15=40y 15/40 =y, 3/8=y

OpenStudy (anonymous):

I'm sorry it was 42 not 40

OpenStudy (anonymous):